OpenStudy (anonymous):

9. Solve -20 = 5/6x X= ?

OpenStudy (alexandervonhumboldt2):

first multply both sides by 6: 5/6x*6=-20*6 5x=-120 now divide both sides by 5: 5x/5=-120/5 x=-24

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

@AlexandervonHumboldt2 can you help me with one more

OpenStudy (alexandervonhumboldt2):

sure

OpenStudy (anonymous):

10. Evaluate x - 4 / 2 when x = -1/6/ -1.6 - 4 / 2 Answer =

OpenStudy (alexandervonhumboldt2):

4/2=2 -1.6-2=-3.6

OpenStudy (anonymous):

Thank you so much